If the tension in a string doubled, then a natural
Answer #1

1) the frequency will increase if the tension is doubled as f \oe sqrt(T)

2) fn = n • f1   so. f1 = f3/3 = 600/3 =200Hz

3) the frequency of the more massive string will be 1/2 of the frequency of the less massive string

as f = (1/2L)sqrt (T/(m/L))

4) increase in the natural frequency as from the above relation, we can see that frequency is inversely proportional to sqrt(L)

5) v = sqrt (T/(m/L)) = sqrt[ 8.5/(0.188/42)] =43.58m/s